OpenStudy (anonymous):

What is the nth term of the arithmetic sequence? 7, 5, 3, 1, … A. –2n B. 7 – 2n C. 9 – 2n D. n – 2

rishavraj (rishavraj):

see the common difference "d" = -2 and first term = 7 \[a _{n} = a _{1} + (n - 1)d\]

rishavraj (rishavraj):

plug the values and get a_n

OpenStudy (anonymous):

I don't know how, what is a? what is n?

rishavraj (rishavraj):

\[a _{n} ~~~is~~~~the~~~~nth~~~term\]

rishavraj (rishavraj):

\[a _{1} ~~~~is~~~the~~~~first~~~term\]

rishavraj (rishavraj):

u r given \[a_1 = 7\]

OpenStudy (anonymous):

yeahhhhhhh still don't get it.

rishavraj (rishavraj):

see the given sequence is 7, 5, 3 , 1, .... the first term i.e \[a_1 = 7\] and the common difference i.e d = -2

OpenStudy (anonymous):

but what is the nth term

rishavraj (rishavraj):

nth term i.e \[a_n \] \[a_n = a_1 + (n - 1)d\] \[a_n = 7 + (n - 1)()\]

rishavraj (rishavraj):

\[a_n = 7 + (n - 1)(-2)\]

OpenStudy (anonymous):

B?

rishavraj (rishavraj):

nah....:(

OpenStudy (anonymous):

A? C? D?

rishavraj (rishavraj):

a_n = a + (n - 1)(-2) a_n = 7 -2n + 2

rishavraj (rishavraj):

a_n = 9 - 2n